OpenStudy (anonymous):

Any help is appreciated(: - A circuit has 12 Ω of resistance and is connected to a 12 v battery. What is the change in power if the resistance decreases to 9 Ω?

OpenStudy (theeric):

Hi! So, I think we can try the technique of finding the power with the \(12\ [\Omega]\) resistor and call it something like \(P_{12}\). Then we can find the power with the \(9\ [\Omega]\) resistor, and call it \(P_9\). Then we can compare the two and find the change in power! You can use any two of voltage, resistance, and current to make a formula for power. So let's look at voltage and resistance. I always remember the simplest equation, \(P=IV\), and then use Ohm's law if I forget the others. Since \(I=\dfrac{V}{R}\), \(P=\dfrac{V^2}{R}\). So now you can calculate \(P_{12}\) and \(P_9\).

OpenStudy (theeric):

The change in power from \(P_{12}\) to \(P_9\) is \(\Delta P=P_9-P_{12}\), always \(\Delta \text{anything}=\text{final}-\text{initial}\).

OpenStudy (anonymous):

Wouldn't that be -3? .-.

OpenStudy (theeric):

What did you get for \(P_9\)?

OpenStudy (anonymous):

I do not understand this at all ._.

OpenStudy (doc.brown):

Less resistance, more stuff can get through. The initial power P=V^2/R P= 12*12/12 P= 12W Change the resistor to a 9ohm P= 12*12/9 P= 16W The change is final minus initial. 16-12 = 4W

OpenStudy (anonymous):

Thanks so much!